OpenStudy (josee):

On what interval(s) is the function f(x) = x^3e^(2x) increasing?

OpenStudy (anonymous):

guess you have to take the derivative and figure out where it is positive (the derivative that is). did you get the derivative?

OpenStudy (anonymous):

\[f'(x)=3x^2e^{2x}+2x^3e^{2x}\] \[=e^{2x}(3x^2+2x^3)\]

OpenStudy (josee):

1) (-inf, -3/2], [0,inf) 2) (-inf, -3/2] 3) (-inf, 0], [3/2, inf) 4) (-inf, 3/2] 5) [-3/2, inf)

OpenStudy (anonymous):

since \[e^{2x}>0\] for all x, you job is to look at \[3x^2+2x^3=x^2(3+2x)\]

OpenStudy (anonymous):

and since \[x^2\geq0\] you only need to solve \[3+2x>0\] \[3>2x\] \[\frac{3}{2}>x\]

OpenStudy (anonymous):

looks like answer 4

OpenStudy (josee):

nopee :/

OpenStudy (anonymous):

yeah because my algebra stinks it is \[3+2x>0\] \[3>-2x\] \[-\frac{3}{2}>x\]

OpenStudy (josee):

lol. i still dont know which interval ;/

OpenStudy (josee):

maye 1 or 2? idk..